OpenStudy (anonymous):

What is the simplified form of the square root of the quantity 98 times n to the nineteenth power.? Choose one answer. a. 2n9the square root of 7. b. 7n4the square root of the quantity 2 times n cubed. c. 7n9the square root of the quantity 2 times n. d. 4n9the square root of the quantity 6 times n.

OpenStudy (anonymous):

Please help me?

OpenStudy (jack1):

sqrt (98n^19) = (98n^19)^(1/2) = (49 x 2 x n^19)^(1/2) = (49)^(1/2) x (2)^(1/2) x (n^19)^(1/2) = 7 x sqrt 2 x sqrt n^19 = 7 x sqrt 2 x n^9.5

OpenStudy (anonymous):

So, C?

OpenStudy (jack1):

C

OpenStudy (anonymous):

Thanks :D

OpenStudy (jack1):

welcome